Need help with number 3. This is for Abstract Algebra Rings and Polynomials
C3 FACTUr COMALETELY Ox4+ 2003+ 3x2+2x+ 2
Answer #1

Answer ) In Q [x], 2t3x+2x+2 = (x+1) (x+2x+2) 2 22z 32422+2 = xt x Z [x, 2t2t3x+22+ 2 = (x) (x) (xti)(X+1) In Z x] x4 2x+3x+2